# Dissociation driven quarkonium spin alignment in Pb–Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV

### v1

@@ I Introduction
The hunt for the deconfined phase of strongly interacting matter known as quark-gluon plasma (QGP) in ultra-relativistic heavy-ion collisions was at its peak, but then something unexpected happened in the grand caves under the Jura mountains. The evidence for QGP formation in heavy-ion collisions has been sought through observables such as collective flow, strangeness enhancement, quarkonia suppression, etc. [1, 2, 3, 4, 5]. Similar to those, a substantial influence of the medium on observables, like strangeness enhancement and collective flow at high multiplicities in ppp-p collisions, was observed at the LHC energies [6, 7, 8]. These findings suggest the possible emergence of QGP-like phenomena in small systems, thereby opening a new research domain to investigate the existence of QGP in such small collision systems. These findings also pose a challenge to established methodologies, particularly for observables that use ppp-p collisions as a baseline, like quarkonia suppression. However, quarkonia, being the bound states of the heavy quark and antiquark pair, have masses significantly larger than the QCD scale, $\Lambda_{\rm QCD}$. The large heavy quark masses originate mainly from their Yukawa couplings to the Higgs field through electroweak symmetry breaking, in contrast to light quarks whose effective masses are largely generated by non-perturbative QCD dynamics and spontaneous chiral symmetry breaking. Therefore, quarkonia, with their unique properties, such as large binding energies, small spatial radii, and distinct sequential melting temperatures, make them a better probe for investigation of deconfined QCD matter  [1]. Since quarkonium states are produced predominantly during the early hard scatterings, the vorticity produced in the earliest times of collisions is expected to affect heavy quarks (anti-quarks) more strongly than light flavors. Through spin-orbit coupling, such early-time vorticity can induce polarization in quarkonium states, which would access the full space–time evolution of the system, and may provide a direct connection between the early dynamics and the final states.
@@
In the absence of any spin alignment, the three spin substates ($m_{j}=+1,0,-1$) are equally probable, giving $\rho_{00}=1/3$. The deviation $\rho_{00}-1/3$ thus quantifies the degree of alignment of the vector meson. A value $\rho_{00}<1/3$ indicates a transverse alignment, whereas $\rho_{00}>1/3$ corresponds to a longitudinal alignment with respect to the chosen quantization axis. A precise determination of $\rho_{00}$ and its associated polarization parameters in different reference frames, such as the helicity, Collins–Soper, and event-plane frames, provides key insights into the spin dynamics of quarkonia and the influence of the surrounding medium in relativistic heavy-ion collisions. However, the recent experimental measurement of spin alignment of $J/\psi$ in Pb–Pb collisions at $\sqrt{s_{\rm NN}}$ = 5.02 TeV is not consistent across different polarization frame of reference. The spin alignment observable $\rho_{00}<$ 1/3 for helicity and event plane frame, while it is observed to be $\rho_{00}>$ 1/3 Collins-Soper frame.
In this study, we investigate the spin alignment of charmonium and bottomonium states, specifically $J/\psi$, $\psi$(2S), $\Upsilon$(1S), and $\Upsilon$(2S) as a function of transverse momentum ($p_{\rm T}$) and charged particle multiplicity in Pb–Pb collisions at $\sqrt{s_{\rm NN}}=5.02~\text{TeV}$. Our study focuses on the spin-dependent dissociation of quarkonia as a possible mechanism for their spin alignment in a vortical QGP medium. Here, we have obtained the spin alignment of quarkonium states considering collisional damping and the gluonic dissociation mechanisms. The combined effect of these in-medium mechanisms on the spin alignment observable $\rho_{00}$ is systematically explored. Further, the space-time evolution of the QGP is modeled using a second-order relativistic viscous hydrodynamic framework, which provides the temperature cooling profile in the expanding medium. The effect of medium vorticity, characterized through the conserved circulation $C$, is further examined to assess its role in generating spin alignment for different quarkonium states within the deconfined phase of thermally equilibrated hot and vortical QCD matter.
@@ II Formulation
The evolution of quarkonium states within the QGP medium is described using the time-independent. Schrödinger equation (SE) and followed by dissociation mechanisms like collisional damping and gluonic dissociation. Collisional damping and gluon-induced dissociation are among the extensively studied processes governing quarkonium suppression in the QGP medium [33, 34, 30, 35, 31, 32]. These mechanisms induce spin-dependent modifications in their dissociation widths, thereby influencing the spin density matrix element, $\rho_{00}$. To quantify this effect, we calculate the spin-dependent decay widths arising from both collisional damping and gluonic dissociation by incorporating the coupling between the spin of the particle and the vorticity of the medium within an effective Hamiltonian framework that includes a medium-modified potential. Given the pivotal role of temperature in modifying the quarkonia characteristics within the QGP medium, it is imperative first to examine the thermodynamic evolution of deconfined QCD matter.
@@ II.1 Thermal Evolution of the Medium
The possible formation of QGP in ultra-relativistic heavy-ion collisions undergoes rapid expansion in both space and time domains due to strong internal pressure gradients. The space–time evolution of such a deconfined medium is proposed to be governed by relativistic hydrodynamics. In the present study, we employ the second-order relativistic viscous hydrodynamics framework to describe the dynamical evolution of the QGP. To solve the hydrodynamic equations, one must specify appropriate initial conditions and an equation of state (EoS) that connects with the thermodynamic quantities of the system. In the absence of a first-principles determination of the initial temperature ($T_{0}$) and thermalization time ($\tau_{0}$) for QGP, we estimate $T_{0}$ using a phenomenological relation [36] that connects it to the experimentally measured charged particle multiplicity:
@@
Multiplicity Class Centrality Class $\frac{dN_{\text{\rm ch}}}{d\eta}>$
@@
Table 1: The $\frac{dN_{\text{\rm ch}}}{d\eta}>$ values measured for Pb+Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV in $|\eta|<$ 0.5 for 11 centrality classes are taken from ALICE experimental data [37].
@@
$\frac{d\phi}{d\tau}=-\frac{2aT\phi}{3b}-\frac{1}{2}\phi\left[\frac{1}{\tau}-\frac{5}{\tau}\frac{dT}{d\tau}\right]+\frac{8aT^{4}}{9\tau}$ (5)
The variable $\phi$ in Eq. 5 is called the shear viscous term, which quantifies the time-dependent change in the shear viscosity ($\eta$). In essence, $\phi$ reflects the characteristics of the medium created under extreme conditions during ultra-relativistic collisions. For the first-order solution of Eq. 4, $\phi$ is defined as $\phi=4\eta/3\tau$. The constants $a$ and $b$ are defined as;
@@
here $N_{f}=3$, is the number of flavors and $\alpha_{s}$ is the strong coupling constant. The first-order viscous correction term, $\phi=4\eta/(3\tau)$, is used to set the initial condition for Eq. 5. By taking the KSS bound of the shear viscosity to entropy density ratio, $\eta/s=1/(4\pi)$, the initial value of the viscous term at $\tau_{0}$ can be expressed as $\phi_{0}=(1/3\pi)(s_{0}/\tau_{0})$. The initial entropy density ($s_{0}$) is estimated using the quasi-particle model (QPM) equation of state at zero baryon chemical potential [30].
@@ II.2 Quarkonia in-medium
Quarkonia traversing through the medium are supposed to be affected by the surrounding partonic environments. The influence of the medium on quarkonia depends on how fast quarkonia is moving in the medium and the relative velocity between the medium and quarkonia. Apart from this, quarkonia, being heavy in nature, do not share the same temperature as the medium. Therefore, to account for the precise impact of the medium on quarkonia, it becomes necessary first to find the effective temperature of the quarkonia within the medium.
@@ II.2.1 Effective Temperature
where, $\theta$ denotes the angle between $v_{m}$ and $v_{Q(nl)}$. The medium temperature $T(\tau,b)$ is obtained from the solution of Eqs. 4 and 5. To obtain an isotropic (angle-independent) representation of the effective temperature, Eq. 8 is averaged over the solid angle, yielding the mean effective temperature as;
@@
The $T_{\rm eff}$ at $T=400$ MeV as a function of $p_{\rm T}$ for different quarkonium states is characterized in Fig. 1. The initial increase in $T_{\rm eff}$ at $p_{\rm T}\leq M_{Q(nl)}$ arises from the effective blue-shift induced when the quarkonium velocity is much smaller than the medium flow velocity, $v_{Q}<<v_{m}$. In this regime, the local thermal distribution sampled by the quarkonium is shifted toward higher apparent temperatures due to the relative motion between the probe and the expanding medium. At $v_{Q}>>v_{m}$, corresponding to the $p_{\rm T}>>M_{Q(nl)}$, the kinematic red-shift dominates, reducing the apparent slope of the thermal spectrum and consequently decreasing $T_{\rm eff}$ for both charmonium and bottomonium states. This systematic behavior indicates that $T_{\rm eff}$ encodes the interplay between quarkonium kinematics and medium flow, and therefore has a direct impact on the in-medium decay width, and eventually influences the spin alignment of quarkonia as it depends on particle decay width (discussed in Sec. II.3).
@@ II.2.2 Quantum characteristic in presence of rotation
For a two-body bound system such as quarkonium immersed in a rotating medium, Eq. 10 can be generalized as;
@@
Here, the index $i=1,2$ refers to the heavy quark and its antiquark, respectively. Applying the standard reduction procedure for a two-body system to an effective one-body problem in the center of the mass frame, the corresponding canonical momenta and position vectors can be defined as;
@@
Taking rotation along the $z$-axis, i.e., $\boldsymbol{\omega}=\omega\hat{z}$, Eq. 14 can be rewritten as;
@@
Using the operators $\mathbf{p}=-i\boldsymbol{\nabla}$, $\mathcal{H}=i\frac{\partial}{\partial t}$, $L_{z}=-i\frac{\partial}{\partial\phi}$, and $S_{1z}+S_{2z}=S_{z}$, we can rewrite the Eq. 15 as;
@@
Using Stokes’ theorem, which relates the line integral to the surface integral of vorticity, Eq. 18 can be rewritten as;
@@
Substituting the expression for $\omega$ in terms of $C$ into Eq. 17 yields the modified radial Schrödinger equation in the rotating frame:
@@
We have numerically solved the Schrödinger equation by incorporating the real part of the medium modified potential, $V(r,m_{D})$, as given by Eq. 21 and obtained the eigen energies and wave functions for the corresponding quarkonium states. Furthermore, using these variables, we quantified the quarkonia dissociation mechanisms in the medium, which are discussed briefly in the following section.
@@ II.2.3 Collisional Damping
The collisional energy loss of quarkonium in the QGP is commonly referred to as collisional damping, which arises due to the interactions of the bound heavy quark pair with the thermal partons in the medium. This effect leads to a finite in-medium width of the quarkonium state, reflecting its partial decoherence and eventual dissociation. The dissociation rate associated with this mechanism can be evaluated using the imaginary part of the potential, $V(r,m_{D})$, which encodes the medium-induced damping effect. The in-medium singlet complex potential for the $c\bar{c}$ bound state is given as [46, 33]:
@@
In Eq. (21), the first two terms on the right-hand side correspond to the screened string and Coulombic interactions, respectively. The third term represents the imaginary part of the potential, which accounts for the collisional damping arising from the exchanged gluons in the QGP medium. Here, $\sigma$ denotes the string tension for the $c\bar{c}$ system, while $m_{D}$ is the Debye screening mass. The strong coupling constant at the hard scale is taken as $\alpha_{s}^{T}$ is obtained under HTL limit, while at the soft scale the effective coupling is defined as $\alpha_{\rm eff}=\tfrac{4}{3}\alpha_{s}^{s}$, where $\alpha_{s}^{s}$ is the coupling evaluated at the relevant momentum-transfer scale for the potential.
@@
$\Gamma_{\rm damp,nl}(\tau,p_{\rm T})=\int[g_{\rm nl}(r)^{\dagger}\left[\rm Im(V)\right]g_{\rm nl}(r)]dr,$ (22)
here $g_{nl}(r)$ is the quarkonium singlet wave function. The spin-dependent wave functions for $J/\psi$, $\psi$(2S), $\Upsilon$(1S) and $\Upsilon$(2S) are obtained by solving the Schrödinger equation (Eq. 20) in the presence of rotation, as described in the preceding section.

### v2

# Dissociation-driven quarkonium spin alignment in Pb–Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV

@@ I Introduction
The hunt for the deconfined phase of strongly interacting matter known as quark-gluon plasma (QGP) in ultra-relativistic heavy-ion collisions was at its peak, but then something unexpected happened in the grand caves under the Jura mountains. The evidence for QGP formation in heavy-ion collisions has been sought through observables such as collective flow, strangeness enhancement, quarkonia suppression, etc. [1, 2, 3, 4, 5]. Similar to those, a substantial influence of the medium on observables, like strangeness enhancement and collective flow at high multiplicities in ppp-p collisions, was observed at the LHC energies [6, 7, 8]. These findings suggest the possible emergence of QGP-like phenomena in small systems, thereby opening a new research domain to investigate the existence of QGP in such small collision systems. These findings also pose a challenge to established methodologies, particularly for observables that use ppp-p collisions as a baseline, like quarkonia suppression. However, quarkonia, being the bound states of the heavy quark and antiquark pair, have masses significantly larger than the QCD scale, $\Lambda_{\rm QCD}$. The large heavy quark masses originate mainly from their Yukawa couplings to the Higgs field through electroweak symmetry breaking, in contrast to light quarks whose effective masses are largely generated by non-perturbative QCD dynamics and spontaneous chiral symmetry breaking. Therefore, quarkonia, with their unique properties, such as large binding energies, small spatial radii, and distinct sequential melting temperatures, make them a better probe for investigation of deconfined QCD matter  [1]. Since quarkonium states are produced predominantly during the early hard scatterings, the vorticity produced in the earliest times of collisions is expected to affect heavy quarks (antiquarks) more strongly than light flavors. Through spin-orbit coupling, such early-time vorticity can induce polarization in quarkonium states, which would access the full space–time evolution of the system, and may provide a direct connection between the early dynamics and the final states.
@@
In the absence of any spin alignment, the three spin substates ($m_{j}=+1,0,-1$) are equally probable, giving $\rho_{00}=1/3$. The deviation $\rho_{00}-1/3$ thus quantifies the degree of alignment of the vector meson. A value $\rho_{00}<1/3$ indicates a transverse alignment, whereas $\rho_{00}>1/3$ corresponds to a longitudinal alignment with respect to the chosen quantization axis. A precise determination of $\rho_{00}$ and its associated polarization parameters in different reference frames, such as the helicity, Collins–Soper, and event-plane frames, provides key insights into the spin dynamics of quarkonia and the influence of the surrounding medium in relativistic heavy-ion collisions. However, the recent experimental measurement of spin alignment of $J/\psi$ in Pb–Pb collisions at $\sqrt{s_{\rm NN}}$ = 5.02 TeV is not consistent across different polarization frame of reference  [28, 29]. The spin alignment observable $\rho_{00}<$ 1/3 for helicity and event plane frame, while it is observed to be $\rho_{00}>$ 1/3 Collins-Soper frame.
In this study, we investigate the spin alignment of charmonium and bottomonium states, specifically $J/\psi$, $\psi$(2S), $\Upsilon$(1S), and $\Upsilon$(2S) as a function of transverse momentum ($p_{\rm T}$) and charged particle multiplicity in Pb–Pb collisions at $\sqrt{s_{\rm NN}}=5.02~\text{TeV}$. This study examines the spin alignment of quarkonia along the vorticity or total angular momentum direction, which is perpendicular to the reaction plane [30]. In the absence of fluctuations, the event plane serves as an estimate of the reaction plane [31]. Therefore, the present calculation is more closely aligned with the event-plane frame. Our study focuses on the spin-dependent dissociation of quarkonia as a possible mechanism for their spin alignment in a vortical QGP medium. Here, we have obtained the spin alignment of quarkonium states considering collisional damping and the gluonic dissociation mechanisms. The combined effect of these in-medium mechanisms on the spin alignment observable $\rho_{00}$ is systematically explored. Further, the space-time evolution of the QGP is modeled using a second-order relativistic viscous hydrodynamic framework, which provides the temperature cooling profile in the expanding medium. The effect of medium vorticity, characterized through the conserved circulation $C$, is further examined to assess its role in generating spin alignment for different quarkonium states within the deconfined phase of thermally equilibrated hot and vortical QCD matter.
@@ II Formulation
The evolution of quarkonium states within the QGP medium is described using the time-independent Schrödinger equation (SE) and followed by dissociation mechanisms like collisional damping and gluonic dissociation. Collisional damping and gluon-induced dissociation are among the extensively studied processes governing quarkonium suppression in the QGP medium [35, 36, 32, 37, 33, 34]. These mechanisms induce spin-dependent modifications in their dissociation widths, thereby influencing the spin density matrix element, $\rho_{00}$. To quantify this effect, we calculate the spin-dependent decay widths arising from both collisional damping and gluonic dissociation by incorporating the coupling between the spin of the particle and the vorticity of the medium within an effective Hamiltonian framework that includes a medium-modified potential. Given the pivotal role of temperature in modifying the quarkonia characteristics within the QGP medium, it is imperative first to examine the thermodynamic evolution of deconfined QCD matter.
@@ II.1 Thermal Evolution of the Medium
The possible formation of QGP in ultra-relativistic heavy-ion collisions undergoes rapid expansion in both space and time domains due to strong internal pressure gradients. The space–time evolution of such a deconfined medium is proposed to be governed by relativistic hydrodynamics. In the present study, we employ the second-order relativistic viscous hydrodynamics framework to describe the dynamical evolution of the QGP. To solve the hydrodynamic equations, one must specify appropriate initial conditions and an equation of state (EoS) that connects with the thermodynamic quantities of the system. In the absence of a first-principles determination of the initial temperature ($T_{0}$) and thermalization time ($\tau_{0}$) of QGP, we estimate $T_{0}$ using a phenomenological relation [Hwa:1984fj] that connects it to the experimentally measured charged particle multiplicity:
@@
Multiplicity Class Centrality Class $\langle\frac{dN_{\text{\rm ch}}}{d\eta}\rangle$
@@
Table 1: The $\langle\frac{dN_{\text{\rm ch}}}{d\eta}\rangle$ values measured for Pb+Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV in $|\eta|<$ 0.5 for 11 centrality classes are taken from ALICE experimental data [39].
@@
$\frac{d\phi}{d\tau}=-\frac{2aT\phi}{3b}-\frac{1}{2}\phi\left[\frac{1}{\tau}-\frac{5}{T}\frac{dT}{d\tau}\right]+\frac{8aT^{4}}{9\tau}$ (5)
The variable $\phi$ in Eq. (5) is called the shear viscous term, which quantifies the time-dependent change in the shear viscosity ($\eta$). In essence, $\phi$ reflects the characteristics of the medium created under extreme conditions during ultra-relativistic collisions. For the first-order solution of Eq. (4), $\phi$ is defined as $\phi=4\eta/3\tau$. The constants $a$ and $b$ are defined as;
@@
here $N_{f}=3$, is the number of flavors and $\alpha_{s}$ is the strong coupling constant. The first-order viscous correction term, $\phi=4\eta/(3\tau)$, is used to set the initial condition for Eq. (5). By taking the KSS bound of the shear viscosity to entropy density ratio, $\eta/s=1/(4\pi)$, the initial value of the viscous term at $\tau_{0}$ can be expressed as $\phi_{0}=(1/3\pi)(s_{0}/\tau_{0})$. The initial entropy density ($s_{0}$) is estimated using the quasi-particle model (QPM) equation of state at zero baryon chemical potential [32].
@@ II.2 Quarkonia in-medium
Quarkonia traversing through the medium are supposed to be affected by the surrounding partonic environments. The influence of the medium on quarkonia depends on how fast quarkonia is moving in the medium and the relative velocity between the medium and quarkonia. Apart from this, quarkonia, being heavy, do not share the same temperature as the medium. Therefore, to account for the precise impact of the medium on quarkonia, it becomes necessary first to find the effective temperature of the quarkonia within the medium.
@@ II.2.1 Effective Temperature
where, $\theta$ denotes the angle between $v_{m}$ and $v_{Q(nl)}$. The medium temperature $T(\tau,b)$ is obtained from the solution of Eqs. (4) and (5). To obtain an isotropic (angle-independent) representation of the effective temperature, Eq. (8) is averaged over the solid angle, yielding the mean effective temperature as;
@@
The $T_{\rm eff}$ at $T=400$ MeV as a function of $p_{\rm T}$ for different quarkonium states is characterized in Fig. 1. The initial increase in $T_{\rm eff}$ at $p_{\rm T}\leq M_{Q(nl)}$ arises from the effective blue-shift induced when the quarkonium velocity is much smaller than the medium flow velocity, $v_{Q}\ll v_{m}$. In this regime, the local thermal distribution sampled by the quarkonium is shifted toward higher apparent temperatures due to the relative motion between the probe and the expanding medium. At $v_{Q}\gg v_{m}$, corresponding to the $p_{\rm T}\gg M_{Q(nl)}$, the kinematic red-shift dominates, reducing the apparent slope of the thermal spectrum and consequently decreasing $T_{\rm eff}$ for both charmonium and bottomonium states. This systematic behavior indicates that $T_{\rm eff}$ encodes the interplay between quarkonium kinematics and medium flow, and therefore has a direct impact on the in-medium decay width, and eventually influences the spin alignment of quarkonia as it depends on particle decay width (discussed in Sec. II.3).
@@ II.2.2 Quantum characteristic in presence of rotation
For a two-body bound system such as quarkonium immersed in a rotating medium, Eq. (10) can be generalized as;
@@
Here, the index $i=1,2$ refers to the heavy quark and its antiquark respectively. Applying the standard reduction procedure for a two-body system to an effective one-body problem in the center of the mass frame, the corresponding canonical momenta and position vectors can be defined as;
@@
Taking rotation along the $z$-axis, i.e., $\boldsymbol{\omega}=\omega\hat{z}$, Eq. (14) can be rewritten as;
@@
Using the operators $\mathbf{p}=-i\boldsymbol{\nabla}$, $\mathcal{H}=i\frac{\partial}{\partial t}$, $L_{z}=-i\frac{\partial}{\partial\phi}$, and $S_{1z}+S_{2z}=S_{z}$, we can rewrite the Eq. (15) as;
@@
Using Stokes’ theorem, which relates the line integral to the surface integral of vorticity, Eq. (18) can be rewritten as;
@@
Substituting the expression for $\omega$ in terms of $C$ into Eq. (17) yields the modified radial Schrödinger equation in the rotating frame:
@@
We have numerically solved the Schrödinger equation by incorporating the real part of the medium modified potential, $V(r,m_{D})$, as given by Eq. (21) and obtained the eigen energies and wave functions for the corresponding quarkonium states. Furthermore, using these variables, we quantified the quarkonia dissociation mechanisms in the medium, which are discussed briefly in the following section.
@@ II.2.3 Collisional Damping
The collisional energy loss of quarkonium in the QGP is commonly referred to as collisional damping, which arises due to the interactions of the bound heavy quark pair with the thermal partons in the medium. This effect leads to a finite in-medium width of the quarkonium state, reflecting its partial decoherence and eventual dissociation. The dissociation rate associated with this mechanism can be evaluated using the imaginary part of the potential, $V(r,m_{D})$, which encodes the medium-induced damping effect. The in-medium singlet complex potential for the heavy quark-antiquark bound state is given as [48, 35]:
@@
In Eq. (21), the first two terms on the right-hand side correspond to the screened string and Coulombic interactions, respectively. The third term represents the imaginary part of the potential, which accounts for the collisional or Landau damping of the low-frequency gauge fields that mediate interactions between quark-antiquark pairs. In Eq. (21), $\sigma$ represents the string tension between the quark-antiquark pair and taken as $\sigma=$ 0.192 GeV2. Further, $m_{D}$ is the Debye screening mass, given as;
$m_{D}=T_{\rm eff}\sqrt{4\pi\alpha_{s}^{T}\left(\frac{N_{c}}{3}+\frac{N_{f}}{6}\right)}$ (22)
here, $N_{c}=3$, $N_{f}=3$ and $\alpha_{s}^{T}$ is the strong coupling constant at the hard scale obtained under the Hard Thermal Loop (HTL) limit. In general, the leading order strong running coupling constant $\alpha_{s}$ is defined as;
$\alpha_{s}(\mu)=\frac{4\pi}{\beta_{0}\ln\left(\frac{\mu^{2}}{\Lambda_{\overline{MS}}^{2}}\right)}$ (23)
where $\beta_{0}=11-\frac{2}{3}N_{f}$, $\Lambda_{\overline{MS}}$ = 0.176 GeV, and $\mu$ is renormalization scale. The $\alpha_{s}^{T}$ is obtained by scaling at the HTL limit ($\mu$ = $2\pi T$), i.e. $\alpha_{s}^{T}=\alpha_{s}(2\pi T)$. As it can be seen from Eq. (23), the $\alpha_{s}^{T}$ depends on medium temperature. At a particular temperature, say T = 400 MeV, $\alpha_{s}^{T}\simeq$ 0.26. Likewise, the coupling constant at the soft scale, $\alpha_{s}^{s}$ and ultra soft scale $\alpha_{s}^{u}$ are determined by evaluating the running coupling at the characteristic scales associated with heavy-quark dynamics; $\alpha_{s}^{s}=\alpha_{s}(m_{q}\alpha_{s}/2)$ and $\alpha_{s}^{u}=\alpha_{s}(m_{q}\alpha_{s}^{2}/2)$. Here, $m_{q}$ is the mass of the charm/bottom quark. Following Eq. (21), effective coupling entering the potential is defined in terms of soft-scale coupling as $\alpha_{\mathrm{eff}}=\frac{4}{3}\alpha_{s}^{s}$. For the chosen set of parameters, we obtain $\alpha_{s}^{s}\simeq 0.33$ for bottom quark, and $\alpha^{s}_{s}\simeq 0.5$ for charm quark, which fixes the numerical value of $\alpha_{\mathrm{eff}}$, respectively.
@@
$\Gamma_{{\rm damp},nlm}(\tau,p_{\rm T})=\int[g_{nlm}(r)^{\dagger}\left[\rm Im(V)\right]g_{nlm}(r)]dr,$ (24)
here $g_{nlm}(r)$ is the quarkonium singlet wave function. The spin-dependent wave functions for $J/\psi$, $\psi$(2S), $\Upsilon$(1S) and $\Upsilon$(2S) are obtained by solving the Schrödinger equation (Eq. (20)) in the presence of rotation, as described in the preceding section.